# Principal Component Analysis and biplots. A Back-to-Basics Comparison of Implementations

### v1

@@ 2.1 A motivating example
Consider the example in Tab.2, and Fig.1. The first two columns of Tab.2 represent the raw data, which are arranged in a so-called data matrix $\mathbf{X}=[x_{ij}]_{n\times m}$ consisting of $i=1,...,n$ observations and $j=1,...,m$ features, or variables. In the example $n=6$ and $m=2$. For each feature $j$ the last two rows of Tab.2 give its mean $\bar{x}_{j}=\frac{1}{n}\sum_{i}x_{ij}$, and sample variance $s_{x_{j}}^{2}=\frac{1}{n-1}\sum_{i}(x_{ij}-\bar{x}_{j})^{2}$. In R this is just: apply(X, 2, mean) and apply(X, 2, var), respectively.
@@ 2.2 Computational aspects of PCA
$\frac{\textrm{d}\mathbf{a}^{T}\mathbf{S}\mathbf{a}}{\textrm{d}\mathbf{a}}-\frac{\textrm{d}}{\textrm{d}\mathbf{a}}\lambda\left(\mathbf{a}^{T}\mathbf{a}-1\right)=\quad$ 00 0
$2\mathbf{a}^{T}\mathbf{S}-2\lambda\mathbf{a}^{T}=\quad$ 00 0
$\left(\mathbf{S}-\lambda\mathbf{I}\right)\mathbf{a}=\quad$ 00 0 (4)
@@ 2.3 Alternative approaches: Singular Value Decomposition
It is worth emphasising that the singular values in $\boldsymbol{\ell}$ are distinct from, but related to the eigenvalues in $\boldsymbol{\lambda}$ (e.g., Hanson and Harvey 2022; Jolliffe 2004, 38). In our example, such relationship is:
@@
 all.equal(abs(unname(U %*D)), abs(unname(Y %*eigen(S)\$vector)))
@@ 2.5 Biplot properties linked to features
[…]
@@
  cos_theta_ab <- (B[1,] %*B[2,]) / (norm(B[1,], "2")*norm(B[2,], "2"))
@@ 3 Discussion of comparative insights
[…]
@@ 3.1 Group 1: SVD
all.equal(as.numeric(test_pca1\$sdev^2), as.numeric(apply(test_pca1\$x,2,var)))
@@
Due to these assumptions, when the argument pc.biplot is (contestably) FALSE the singular values returned by biplot.prcomp are close to, but different from $\boldsymbol{\ell}$ from Equation 2.3:
@@ 3.2 Group 2: eigenvalue problem
A natural starting point is the function princomp() in base-R. Nearly a homonym of prcomp(), this implementation relies upon the eigenvalues and eigenvectors of the covariance matrix $\mathbf{S}$, rather than the outputs of a data-matrix SVD. Yet princomp() disagrees with Equation 2.2 in that it uses the number of observations $n$, not $n-1$, as the divisor for the covariance matrix. What is more, the signs of the scores returned by princomp() differ from those computed directly using eig() due to further processing within the function:
@@
test_pca5_eig <- eigen(1/nrow(Y)*t(Y) %*Y, symmetric = TRUE)\$vectors
all.equal(abs(unname(Y %*test_pca5_eig)), abs(unname(test_pca5\$scores)))
@@
 all.equal(as.numeric(test_pca5\$sdev^2), as.numeric(apply(test_pca5\$scores,2,var)))
@@
test_pca7_eig <- eigen(t(Y) %*Y, symmetric = FALSE)
@@ 3.3 Group 3: generalised SVD
all.equal(test_pca9\$ind\$coord, test_pca9\$svd\$%*D)
@@
all.equal(test_pca9\$eig, test_pca9\$svd\$vs^2)
@@
all.equal(test_pca9\$eig, unname(eigen(1/nrow(Y)*t(Y) %*Y)\$values))
@@
all.equal(test_pca9\$var\$coord, test_pca9\$svd\$%*D)

### v2

# A peer reviewed, substantially rewritten version of this preprint is available here and freely accessible:

@@ 2.1 A motivating example
Consider the example in Tab.2, and Fig.1. The first two columns of Tab.2 represent the raw data, which are arranged in a so-called data matrix $\mathbf{X}=[x_{ij}]_{n\times m}$ consisting of i=1,,ni=1,...,n observations and j=1,,mj=1,...,m features, or variables. In the example $n=6$ and $m=2$. For each feature $j$ the last two rows of Tab.2 give its mean $\bar{x}_{j}=\frac{1}{n}\sum_{i}x_{ij}$, and sample variance $s_{x_{j}}^{2}=\frac{1}{n-1}\sum_{i}(x_{ij}-\bar{x}_{j})^{2}$. In R this is just: apply(X, 2, mean) and apply(X, 2, var), respectively.
@@ 2.2 Computational aspects of PCA
$\frac{\textrm{d}\mathbf{a}^{T}\mathbf{S}\mathbf{a}}{\textrm{d}\mathbf{a}}-\frac{\textrm{d}}{\textrm{d}\mathbf{a}}\lambda\left(\mathbf{a}^{T}\mathbf{a}-1\right)=\quad$ $0$
$2\mathbf{a}^{T}\mathbf{S}-2\lambda\mathbf{a}^{T}=\quad$ $0$
$\left(\mathbf{S}-\lambda\mathbf{I}\right)\mathbf{a}=\quad$ $0$ (4)
@@ 2.3 Alternative approaches: Singular Value Decomposition
[…]
@@
 all.equal(abs(unname(U %*% D)), abs(unname(Y %*% eigen(S)\$vector)))
@@ 2.5 Biplot properties linked to features
which is the definition of cosine similarity. In Equation 15 the angle in radians between two features vectors is $\theta$; and $\langle{\cdot,\cdot}\rangle$ denotes the inner product between them. What is rarely shown is that Equation 15 is equivalent to the correlation between the two features in our simple example. This equivalence can be proved starting from the correlation coefficient between two columns of matrix $\mathbf{Y}$:
@@
  cos_theta_ab <- (B[1,] %*% B[2,]) / (norm(B[1,], "2")*norm(B[2,], "2"))
@@ 3 Discussion of comparative insights
   FactoMineR PCA() a svd\$V $\frac{\ell_{j}}{\sqrt{n}}=\texttt{svd\$vs}$ ind\$coord $\circ$
   PCAmixdata PCAmix() a svd\$V $\frac{\ell_{j}}{\sqrt{n}}=\texttt{svd\$vs}$ ind\$coord $\circ$
@@ 3.1 Group 1: SVD
all.equal(as.numeric(test_pca1\$sdevˆ2), as.numeric(apply(test_pca1\$x,2,var)))
@@
Due to these assumptions, when the argument pc.biplot is (contestably) FALSE the singular values returned by biplot.prcomp are close to, but different $\boldsymbol{\ell} from$ Equation 2.3:
@@ 3.2 Group 2: eigenvalue problem
A natural starting point is the function princomp() in base-R. Nearly a homonym of prcomp(), this implementation relies upon the eigenvalues and eigenvectors of the covariance matrix $\mathbf{S}$, rather than the outputs of a data-matrix SVD. Yet princomp() disagrees with Equation 2.2 in that it uses the number of observations $n$, not n1n-1, as the divisor for the covariance matrix. What is more, the signs of the scores returned by princomp() differ from those computed directly using eig() due to further processing within the function:
@@
test_pca5_eig <- eigen(1/nrow(Y)*t(Y) %*% Y, symmetric = TRUE)\$vectors
all.equal(abs(unname(Y %*% test_pca5_eig)), abs(unname(test_pca5\$scores)))
@@
 all.equal(as.numeric(test_pca5\$sdevˆ2), as.numeric(apply(test_pca5\$scores,2,var)))
@@
test_pca7_eig <- eigen(t(Y) %*% Y, symmetric = FALSE)
@@ 3.3 Group 3: generalised SVD
all.equal(test_pca9\$ind\$coord, test_pca9\$svd\$%*% D)
@@
all.equal(test_pca9\$eig, test_pca9\$svd\$vsˆ2)
@@
all.equal(test_pca9\$eig, unname(eigen(1/nrow(Y)*t(Y) %*% Y)\$values))
@@
all.equal(test_pca9\$var\$coord, test_pca9\$svd\$%*% D)